# Distillation Robustifies Unlearning

### v2

@@ Abstract
Current LLM unlearning methods are not robust: they can be reverted easily with a few steps of finetuning. This is true even for the idealized unlearning method of training to imitate an oracle model that was never exposed to unwanted information, suggesting that output-based finetuning is insufficient to achieve robust unlearning. In a similar vein, we find that training a randomly initialized student to imitate an unlearned model transfers desired behaviors while leaving undesired capabilities behind. In other words, distillation robustifies unlearning. Building on this insight, we propose Unlearn-Noise-Distill-on-Outputs (UNDO), a scalable method that distills an unlearned model into a partially noised copy of itself. UNDO introduces a tunable tradeoff between compute cost and robustness, establishing a new Pareto frontier on synthetic language and arithmetic tasks. At its strongest setting, UNDO matches the robustness of a model retrained from scratch with perfect data filtering while using only 60-80% of the compute and requiring only 0.01% of the pretraining data to be labeled. We also show that UNDO robustifies unlearning on the more realistic Weapons of Mass Destruction Proxy (WMDP) benchmark. Since distillation is widely used in practice, incorporating an unlearning step beforehand offers a convenient path to robust capability removal.
@@ 1 Introduction
This finding has immediate practical implications. Distillation is already standard practice in many contexts [29, 82]. For example, LLM developers distill models to reduce inference costs [71, 72]. By applying unlearning methods before distillation, these developers can produce models that are strongly resistant to relearning attacks.
@@
  • In Section 3, we provide evidence of a limitation in behavioral training for robust unlearning in LLMs. Even when achieving perfect behavioral agreement with an unlearning oracle [24], models retain latent capabilities that can be easily restored through finetuning.
  • In Section 4, we show that distillation robustifies unlearning. Training a new, randomly initialized model on the outputs of an unlearned model will transfer desired behavior while leaving latent capabilities behind.
  • In Section 5, we propose UNDO, which distills an unlearned model’s outputs into a noised copy of itself. This method enables a novel compute-robustness tradeoff that interpolates between mere behavioral suppression and fully robust unlearning.
@@ 2 Robust Unlearning
Elicitation Mechanism. We use finetuning on the forget set, commonly referred to as a relearning attack, as our elicitation method. Relearning attacks have repeatedly proven to be among the most effective ways to elicit unlearned knowledge or capabilities in supposedly unlearned models [31, 58]. Therefore, recent studies treat resistance to such finetuning-based probes as the primary criterion for robust unlearning [11, 22, 16]. For rigorous evaluation, we apply several relearning attacks at different learning rates and measure the maximum elicited forget performance of all attacks.
@@ 3 Oracle Matching Does Not Guarantee Robust Unlearning
Unlearning methods often finetune the outputs of the reference model. For example, Gradient Ascent combined with Gradient Descent maximizes cross-entropy loss on the forget set and minimizes it on the retain set [48]. The aspirational goal of such procedures is oracle matching: attaining outputs indistinguishable from those of a model that never encountered the undesired data.
@@
Datasets and Models. We conduct experiments in a language setting and an arithmetic setting, defined below. For each setting, we train a reference model with both retain and forget capabilities, and an oracle model with only retain capabilities by construction.
More details on how these datasets and reference models were generated can be found in Appendix A.
  • Language. The retain set is English documents from FineWebEdu [56]; the forget set is Korean documents from FineWeb2 [57]; the training data is the union of these two sets. The reference model is a 100M parameter model based on the Gemma 2 architecture (see Table 1 for details) pretrained on 2B tokens, 1B from retain data and 1B from forget; the oracle model is the same 100M model architecture pretrained on 1B tokens of the retain set. The retain and forget evaluations are cross-entropy loss on held-out portions of the retain and forget sets, respectively, each containing 500K tokens.
@@
Oracle Matching Training. We match the reference model’s outputs with those of the oracle model using the Kullback-Leibler Divergence (KL Divergence) as the loss function. For comparison, we also match a randomly initialized model to the oracle model in the same way. The hyperparameters of the distillation setup can be found in Table 3 and the loss curves are shown in Figure 2 (a), where we use the label Student (Reference) or Student (Random) depending on whether the student in the distillation setup is the reference model or the randomly initialized model, respectively. In both settings, the distillation loss tends to zero.
Relearning. After distillation, we apply relearning attacks to all three models, Student (Reference), Student (Random), and Oracle Teacher, for 500 steps on data from the forget set. The hyperparameters for this training are in Table 4. We show the corresponding learning and accuracy curves for multiple learning rates in Figure 2 (b) and (c). The bolded lines correspond to the worst-case adversary (maximum elicitation) for that number of training steps. In both the arithmetic and language settings, Student (Reference) quickly learns the forget distribution compared to the Student (Random) and Oracle Teacher models, which are more robust. This pattern holds both for the worst-case adversary and uniformly across all learning rates.
@@
As illustrated in Figure 2, distilling the oracle’s outputs into a randomly initialized student (Student (Random)) results in a model whose relearning speed on the forget distribution closely matches that of the oracle itself. Given that existing machine unlearning methods provide practical means of approximating oracle outputs, an important next question is whether this robustness to relearning remains after replacing true oracle outputs with approximations. In the next section, we demonstrate empirically that they do.
@@ 4 Distillation Robustifies Unlearning
Unlearning Method C: Representation Misdirection for Unlearning (RMU) operates at the representation level rather than the output level. It minimizes $\mathcal{L}(\theta)\coloneq L_{\text{misdirect}}(\theta)+L_{\text{preserve}}(\theta)$, where $L_{\text{misdirect}}$ uses mean squared error (MSE) loss to push the model’s internal activations at specific layers toward a random direction for forget examples, and $L_{\text{preserve}}$ applies MSE to maintain similarity to the original model’s activations for retain examples.
@@
While RMU combined with distillation shows less impressive results in the arithmetic domain in Figure 4 (f), it still offers marked improvement over the unlearning-only model. We observe that such underperformance can occur when the initial retain-forget trade-off established by the unlearning method is less favorable. In Figure 3, RMU achieved only 62% retain and 6.8% forget accuracy, compared to MaxEnt’s superior 80% and 1.3% (Section B.1). This suggests that distillation effectiveness can depend on the quality of the teacher model’s retain-forget tradeoff.
@@
These results provide empirical evidence validating claims of prior work about the challenges of robust unlearning. Some have argued that quick relearning of supposedly unlearned capabilities stems from latent information preserved in the parameter space [88, 55, 64] due to properties like plasticity [45] and representational entanglement [89, 48]. Our experiments confirm that when we apply standard unlearning methods to a model, the resulting model relearns quickly under finetuning. However, when we distill this unlearned model into a randomly initialized student, the student relearns substantially slower while maintaining the same unlearning behavior.
@@ 5 Trading Compute for Unlearning Robustness
So far, we have demonstrated that distilling an unlearned model into a randomly initialized model significantly enhances robustness against relearning attacks. However, this approach introduces substantial computational costs, as we are training a student model from scratch. Moving beyond the basic Unlearn-and-Distill method established in Section 4, we now investigate whether the method can be adapted to achieve different levels of robustness with different amounts of compute. That is, can we trade off a bit of robustness to reduce compute costs, or vice versa?
In this section, we conduct experiments that generalize the Unlearn-and-Distill approach in Section 4 with a three-step process. (i) Unlearn: create a behaviorally suppressed model (teacher) by applying standard unlearning methods. (ii) Noise: create a student model by perturbing the weights of the suppressed model, damaging it. (iii) Distill: repair this damaged student by distilling to recover the teacher’s original behavior. The controlled perturbation enables us to interpolate between the original parameters and full reinitialization. We refer to this method as Unlearn-Noise-Distill-on-Outputs, or simply, UNDO. We define the perturbed model as:
@@
where mixing coefficient $\alpha\in[0,1]$, noise scale $\beta\in\mathbb{R}^{+}$, and $N$ represents sampled noise. We sample $N$ using Xavier initialization, though other random noise sampling methods could be explored. Intuitively, the term $(1-\alpha)\theta_{\text{suppressed}}$ shrinks [2] the suppressed model’s parameters, and the term $\alpha\beta N$ injects noise. As we vary $\alpha$ from 0 to 1, we can view $\theta_{\text{perturbed}}$ as interpolating between the original paramterization and full reinitialization. We observe later in Section 6 that simply shrinking the parameters (setting $\beta=0$) can also be effective, as the key idea is to globally damage the network to varying degrees, though including $\beta$ adds expressivity to the formulation. Note that when $\alpha=1$ and $\beta=1$, the perturbation reduces to random initialization.
Experimental Setup. We start by perturbing the models that are suppressed with MaxEnt from Section 4 and follow the same protocols for distillation using forward KL divergence on the outputs. For both language and arithmetic settings, we experiment with various values [^1] We test $\alpha\in\{0.1,0.2,0.3,0.4,0.5,0.55,0.6,0.65,0.7,0.75,0.8\}$ for language and arithmetic tasks. of $\alpha$ while fixing $\beta=0.1$. For each perturbation level, we perform distillation until the model reaches at least 95% of the teacher model’s performance on the retain set or completes training on one full epoch of pretraining data, whichever comes first. Lower-level experimental details are available in Appendix C.
Figure 5: Unlearning robustness scales with more perturbation. (a, c) UNDO scaling trend for $\alpha$ between 0.1 and 0.8 and $\beta=0.1$, showing trade-off between robustness measured as $(P_{\text{UNDO}}-P_{\text{Unlearn Only}})/(P_{\text{Data Filtering}}-P_{\text{Unlearn Only}})$ where $P$ is performance, and compute measured as $S_{\text{UNDO}}/S_{\text{Data Filtering}}$ where $S$ is training steps. Points denote median values, error bars show variation across five random seeds. (b) Relearning trends for Korean domain with $\alpha=\{0.2,0.4,0.6,0.8\}$. (d) Relearning trends for Multiplication & Division with $\alpha=\{0.55,0.65,0.7,0.75\}$.
@@
To learn more about the interplay between mixing coefficient ($\alpha$) and noise scale ($\beta$), we conduct a hyperparameter sweep as reported in Figure 6. Changes in mixing coefficient ($\alpha$) produce a more pronounced effect on robustness than the noise scale ($\beta$), which has a less significant impact. This pattern holds across all settings where models are distilled to within ±$\pm 5\%$ of the teacher’s retain performance. While incorporating standard noise from Xavier initialization can help improve robustness for very low $\alpha$ values ($\alpha=0.1$), we find that for most $\alpha$ values tested, the specific value of $\beta$ had less impact. Interestingly, non-standard $\beta$ values (0.05, 0.1, or 0.5) typically worked similarly well or better than $\beta=1.0$ in our language setup. The computational cost matrix in Figure 6 (e) confirms that $\alpha$ is the primary driver of compute requirements, with higher values often requiring more distillation steps.
@@ 6 Comparisons with Other Unlearning Methods
We compare against several approaches that address the robust unlearning problem using different mechanisms to resist relearning attacks. Sharpness-Aware Minimization (SAM) [22] optimizes for flat loss regions where small parameter perturbations maintain similar outputs, making the model less sensitive to fine-tuning. Representation Noising (RepNoise) [59] combines unlearning with noise injection at the representation level, creating interference in the model’s internal representations of forget capabilities. UNDIAL [19] also uses distillation for unlearning but differs from our approach by not applying global parameter corruption techniques like noising. For UNDO, we fix $\alpha$ to 0.6 and $\beta$ to 0.1, and distill from a MaxEnt-suppressed model with varying retain thresholds [^2] We test {0.05, 0.09, 0.13, 0.17, 0.21, 0.25, 0.29, 0.33} retrain thresholds for language and arithmetic tasks. to obtain different points along the retain-forget trade-off curve. Figure 5 suggests that higher $\alpha$ values would yield greater robustness, while lower $\alpha$ values would yield less. We discuss other relevant experimental details in Appendix D.
@@
Moving beyond our synthetic language and arithmetic unlearning tasks, we also test our methods on the more realistic Weapons of Mass Destruction Proxy (WMDP) benchmark [40]. For this setup, we use Gemma-2-2B [71] rather than custom pretrained models. Following our established methodology, we first apply standard unlearning using either MaxEnt or RMU, then apply UNDO with $\alpha=0.25$ and $\beta=0$. We compare this approach against the other robust unlearning baselines. We evaluate robustness against seven diverse adversarial scenarios involving different data mixtures, question-answer formats, model perturbations, and learning rate variations (See Section D.2).
@@
Limitations. In our WMDP experiments on a pretrained Gemma-2-2B model, the UNDO model achieves 4.88% lower on MMLU on average. While this places UNDO on the Pareto frontier, other methods achieve a similar trade-off between robust forget and retain performance in this setting, as shown in Figure 11. This contrasts with the prior arithmetic and language setting results, where we see clear gains on the Pareto frontier in Figure 7. We hypothesize this difference comes from the fact that we undertrain in the WMDP setting, distilling on only 0.015% of the size of the original pretraining corpus compared to higher percentages in the other synthetic settings. We discuss this more in Section D.2.7). We expect that model developers would not encounter the same challenges and would have sufficient access to compute, pretraining data, and their proprietary methods to accelerate the process. Distillation is considerably more costly than finetuning-based unlearning methods, but we expect this cost to be justified when robust capability removal is critical.

### v3

@@ Abstract
Current LLM unlearning methods are not robust. A few steps of finetuning can revert their effects. We begin by showing that this is true even for an idealized form of unlearning: training to imitate a model that was never trained on unwanted information. This shows that training a model can drastically modify its input-output behavior while leaving its underlying capabilities intact. In light of this dynamic, we show our main result. Training a randomly initialized student on the outputs of an unlearned model transfers behaviors while leaving latent capabilities behind. In short, distillation robustifies unlearning. Based on this result, we propose Unlearn-Noise-Distill-on-Outputs (UNDO), a scalable method that distills an unlearned model into a noised copy of itself. UNDO introduces a tunable tradeoff between compute cost and robustness, establishing a new Pareto frontier on synthetic language and arithmetic tasks. At its strongest setting, UNDO matches the robustness of a model retrained from scratch with perfect data filtering while using only 60-80% of the compute and requiring only 0.01% of the pretraining data to be labeled. We also show that UNDO robustifies unlearning on the more realistic Weapons of Mass Destruction Proxy (WMDP) benchmark. Since distillation is widely used in practice, incorporating an unlearning step beforehand offers a convenient path to robust capability removal.
@@ 1 Introduction
This finding has immediate practical implications. Distillation is commonly used in frontier AI development [30, 86]. For example, LLM developers distill models to reduce inference costs [74, 75]. By applying unlearning methods before distillation, these developers can produce models that are strongly resistant to relearning attacks.
@@
  • In Section˜3, we provide evidence of a limitation in behavioral training for robust unlearning in LLMs. Even when achieving perfect behavioral agreement with an unlearning oracle [25], models retain latent capabilities that can be easily restored through finetuning.
  • In Section˜4, we show that distillation robustifies unlearning. Training a new, randomly initialized model on the outputs of an unlearned model will transfer desired behavior while leaving latent capabilities behind.
  • In Section˜5, we propose UNDO, which distills an unlearned model’s outputs into a noised copy of itself. This method enables a novel compute-robustness tradeoff that interpolates between mere behavioral suppression and fully robust unlearning.
@@ 2 Robust Unlearning
Elicitation Mechanism. We use finetuning on the forget set, commonly referred to as a relearning attack, as our elicitation method. Relearning attacks have repeatedly proven to be among the most effective ways to elicit unlearned knowledge or capabilities in supposedly unlearned models [31, 33, 60]. Therefore, recent studies treat resistance to such finetuning-based probes as the primary criterion for robust unlearning [11, 23, 16]. For rigorous evaluation, we apply several relearning attacks at different learning rates and measure the maximum elicited forget performance of all attacks.
@@ 3 Oracle Matching Does Not Guarantee Robust Unlearning
Unlearning methods often finetune the outputs of the reference model. For example, Gradient Ascent combined with Gradient Descent maximizes cross-entropy loss on the forget set and minimizes it on the retain set [50]. The aspiration of such procedures is oracle matching: attaining outputs indistinguishable from those of a model that never encountered the undesired data.
@@
Datasets and Models. We conduct experiments in a language setting and an arithmetic setting, defined below. For each setting, we train a reference model with both retain and forget capabilities, and an oracle model with only retain capabilities by construction. More details on how these datasets and reference models were generated can be found in Appendix A.
  • Language. The retain set is English documents from FineWebEdu [58]; the forget set is Korean documents from FineWeb2 [59]; the training data is the union of these two sets. The reference model is a 100M parameter model based on the Gemma 2 architecture (see Table˜1 for details) pretrained on 2B tokens, 1B from retain data and 1B from forget; the oracle model is the same 100M model architecture pretrained on 1B tokens of the retain set. The retain and forget evaluations are cross-entropy loss on held-out portions of the retain and forget sets, respectively, each containing 500K tokens.
@@
Oracle Matching Training. We match the reference model’s outputs with those of the oracle model using the Kullback-Leibler Divergence (KL Divergence) as the loss function. For comparison, we also match a randomly initialized model to the oracle model in the same way. The hyperparameters of the distillation setup can be found in Table˜3 and the loss curves are shown in Figure˜2 (a), where we use the label Student (Reference) or Student (Random) depending on whether the student in the distillation setup is the reference model or the randomly initialized model, respectively. In both settings, the distillation loss tends to zero.
Relearning. After distillation, we apply relearning attacks to all three models, Student (Reference), Student (Random), and Oracle Teacher, for 500 steps on data from the forget set. The hyperparameters for this training are in Table˜4. We show the corresponding learning and accuracy curves for multiple learning rates in Figure˜2 (b) and (c). The bolded lines correspond to the worst-case adversary (maximum elicitation across learning rates) for that number of training steps. In both the arithmetic and language settings, Student (Reference) quickly learns the forget distribution compared to the Student (Random) and Oracle Teacher models, which are more robust. This pattern holds both for the worst-case adversary and uniformly across all learning rates.
@@
As illustrated in Figure˜2, distilling the oracle’s outputs into a randomly initialized student (Student (Random)) results in a model whose relearning speed on the forget distribution closely matches that of the oracle itself. Given that existing machine unlearning methods provide practical means of approximating oracle outputs, an important next question is whether this robustness to relearning remains after replacing true oracle outputs with approximations. In the next section, we demonstrate empirically that they do.
@@ 4 Distillation Robustifies Unlearning
Unlearning Method C: Representation Misdirection for Unlearning (RMU) operates at the representation level rather than the output level. It minimizes $\mathcal{L}(\theta)\coloneq L_{\text{misdirect}}(\theta)+\alpha\cdot L_{\text{preserve}}(\theta)$, where $L_{\text{misdirect}}$ uses mean squared error (MSE) loss to push the model’s internal activations at specific layers toward a random direction for forget examples, and $L_{\text{preserve}}$ applies MSE to maintain similarity to the original model’s activations for retain examples.
@@
While RMU combined with distillation shows less impressive results in the arithmetic domain in Figure 4 (f), it still offers marked improvement over the unlearning-only model. We observe that such underperformance can occur when the initial retain-forget trade-off established by the unlearning method is less favorable. In Figure 3, RMU achieved only 62% retain and 6.8% forget accuracy, compared to MaxEnt’s superior 80% and 1.3% (Appendix B.1). This suggests that distillation effectiveness can depend on the quality of the teacher model’s retain-forget tradeoff.
@@
These results provide some empirical evidence supporting claims of prior work about the challenges of robust unlearning. They have argued that quick relearning of supposedly unlearned capabilities stems from latent information preserved in the parameter space [92, 57, 66] due to properties like plasticity [47] and representational entanglement [93, 50]. Our experiments confirm that when we apply standard unlearning methods to a model, the resulting model relearns quickly under finetuning. However, when we distill this unlearned model into a randomly initialized student, the student relearns substantially slower as there was no latent capability to begin with.
Extension to Post-Training Unlearning. While our main experiments distill into randomly initialized networks to remove capabilities acquired during pretraining, this approach also applies when undesired behavior is learned during post-training. In such cases, the original base model (which never possessed the undesired capability) can serve as the distillation target. In Appendix F, we demonstrate this on the TOFU benchmark [50].
@@ 5 Trading Compute for Unlearning Robustness
So far, we have demonstrated that distilling an unlearned model into a randomly initialized model significantly enhances robustness against relearning attacks. However, this approach introduces substantial computational costs, as we are training a student model from scratch. Moving beyond the basic Unlearn-and-Distill method established in Section˜4, we now investigate whether the method can be adapted to achieve different levels of robustness with different amounts of compute. That is, can we trade off robustness to reduce compute costs.
In this section, we conduct experiments that generalize the Unlearn-and-Distill approach in Section˜4 with a three-step process. (i) Unlearn: create a behaviorally suppressed model (teacher) by applying standard unlearning methods. (ii) Noise: create a student model by perturbing the weights of the suppressed model, damaging it. (iii) Distill: repair this damaged student by distilling to recover the teacher’s original behavior. The controlled perturbation enables us to interpolate between the original parameters and full reinitialization. We refer to this method as Unlearn-Noise-Distill-on-Outputs, or simply, UNDO. We define the perturbed model as:
@@
where mixing coefficient $\alpha\in[0,1]$, noise scale $\beta\in\mathbb{R}^{+}$, and $N$ represents sampled noise. We sample $N$ using Xavier initialization, though other random noise sampling methods could be explored. Intuitively, the term $(1-\alpha)\theta_{\text{suppressed}}$ shrinks [2] the suppressed model’s parameters, and the term $\alpha\beta N$ injects noise. As we vary $\alpha$ from 0 to 1, we can view $\theta_{\text{perturbed}}$ as interpolating between the original paramterization and full reinitialization. We observe later in Section˜6 that simply shrinking the parameters (setting $\beta=0$) can also be effective, as the key idea is to globally damage the network to varying degrees, though including $\beta$ adds expressivity to the formulation. Note that when $\alpha=1$ and $\beta=1$, the perturbation reduces to random initialization.
Experimental Setup. We start by perturbing the models that are suppressed with MaxEnt from Section˜4 and follow the same protocols for distillation using forward KL divergence on the outputs. For both language and arithmetic settings, we experiment with various values [^1] We test $\alpha\in\{0.1,0.2,0.3,0.4,0.5,0.55,0.6,0.65,0.7,0.75,0.8\}$ for language and arithmetic tasks. of $\alpha$ while fixing $\beta=0.1$. For each perturbation level, we perform distillation until the model reaches at least 95% of the teacher model’s performance on the retain set or completes training on one full epoch of pretraining data, whichever comes first. Low-level experimental details are available in Appendix C.
Figure 5: Unlearning robustness scales with more perturbation. (a, c) UNDO scaling trend for $\alpha$ between 0.1 and 0.8 and $\beta=0.1$, showing trade-off between robustness measured as $(P_{\text{UNDO}}-P_{\text{Unlearn Only}})/(P_{\text{Data Filtering}}-P_{\text{Unlearn Only}})$ where $P$ is forget performance, and compute measured as $S_{\text{UNDO}}/S_{\text{Data Filtering}}$ where $S$ is training steps. Points denote median values, error bars show variation across five random seeds. (b) Relearning trends for Korean domain with $\alpha=\{0.2,0.4,0.6,0.8\}$. (d) Relearning trends for Multiplication & Division with $\alpha=\{0.55,0.65,0.7,0.75\}$.
@@
To learn more about the interplay between mixing coefficient ($\alpha$) and noise scale ($\beta$), we conduct a hyperparameter sweep as reported in Figure 6. Changes in mixing coefficient ($\alpha$) produce a more pronounced effect on robustness than the noise scale ($\beta$), which has a less significant impact. This pattern holds across all settings where models are distilled to within $\pm 5\%$ of the teacher’s retain performance. While incorporating standard noise from Xavier initialization can help improve robustness for very low $\alpha$ values ($\alpha=0.1$), we find that for most $\alpha$ values tested, the specific value of $\beta$ had less impact. Interestingly, we observe that using lower $\beta$ values (0.05, 0.1, or 0.5) typically worked similarly well or better than $\beta=1.0$ in our language setup. The computational cost matrix in Figure 6 (e) confirms that $\alpha$ is the primary driver of compute requirements, with higher values often requiring more distillation steps.
@@ 6 Comparisons with Other Unlearning Methods
We also compare against several approaches that address the robust unlearning problem using different mechanisms to resist relearning attacks. Sharpness-Aware Minimization (SAM) [23] optimizes for flat loss regions where small parameter perturbations maintain similar outputs, making the model less sensitive to finetuning. Representation Noising (RepNoise) [61] combines unlearning with noise injection at the representation level, creating interference in the model’s internal representations of forget capabilities. UNDIAL [19] also uses distillation for unlearning but differs from our approach by not applying global parameter corruption techniques like noising. For UNDO, we fix $\alpha$ to 0.6 and $\beta$ to 0.1, and distill from a MaxEnt-suppressed model with varying retain thresholds [^2] We test {0.05, 0.09, 0.13, 0.17, 0.21, 0.25, 0.29, 0.33} retrain thresholds for language and arithmetic tasks. to obtain different points along the retain-forget trade-off curve. Figure 5 suggests that higher $\alpha$ values would yield greater robustness, while lower $\alpha$ values would yield less. We discuss other relevant experimental details in Appendix D.
@@
Moving beyond our synthetic language and arithmetic unlearning tasks, we also test our methods on the more realistic Weapons of Mass Destruction Proxy (WMDP) benchmark [42]. For this setup, we use Gemma-2-2B [74] rather than custom pretrained models. Following our established methodology, we first apply standard unlearning using either MaxEnt or RMU, then apply UNDO with $\alpha=0.25$ and $\beta=0$. We compare this approach against the other robust unlearning baselines. We evaluate robustness against seven diverse adversarial scenarios involving different data mixtures, question-answer formats, model perturbations, and learning rate variations (See Appendix D.2).
@@
Limitations. In our WMDP experiments on a pretrained Gemma-2-2B model, the UNDO model achieves 4.88% lower on MMLU on average. While this places UNDO on the Pareto frontier, other methods achieve a similar trade-off between robust forget and retain performance in this setting, as shown in Figure 11. This contrasts with the prior arithmetic and language setting results, where we see clear gains on the Pareto frontier in Figure 7. We hypothesize this difference comes from the fact that we undertrain in the WMDP setting, distilling on only 0.015% of the size of the original pretraining corpus compared to higher percentages in the other synthetic settings. We discuss this more in Appendix D.2.7. We expect that model developers would not encounter the same challenges and would have sufficient access to compute, pretraining data, and their proprietary methods to accelerate the process. Distillation is considerably more costly than finetuning-based unlearning methods, but we expect this cost to be justified when robust capability removal is critical.